of this year's graduating seniors at south high, 9/10 will be going to college. Of these, 4/5 will go to four year college, while the rest will be going to two- year colleges. What part of the class will be going to two-year colleges?: of this year's graduating seniors at south high, 9/10 will be going to college. Of these, 4/5 will go to four year college, while the rest will be going to two- year colleges. What part of the class will be going to two-year colleges?
Answer by leo4frio(2)

Wow, 90% of the students go on to college, what a prestigious high school! A key phrase in this problem is "Of these"; it lets us know that we're looking for a part of a part, a fraction of a fraction. In other words, we'll be multiplying two fractions, but make sure you realize what fractions need to be multiplied. Is it 9/10 and 4/5? Nope, that would tell us what part of South High is going to a four year school. Of those going on to college, 4/5 are going to four year colleges so 1/5 must be going on to a two year college (5/5 - 4/5 = 1/5). So, to find the portion or part of South High that's not only going on to college but specifically to a two year college we must multiply 9/10 by 1/5. Since 9 x 1 = 9 and 10 x 5 = 50, our answer is 9/50. Notice: 9 and 50 don't have any common factors, so our answer stands at 9/50 because this fraction can not be simplified.